chasity and megan are riding their bicycles to raise money for the Catholic Charities childrens home. they are 45 miles apart when they begin cycling towards one another and they both start cycling at the same time.if Chasity travels at 14 miles per hour and Megan travels 16 miles per hour, how long will take them to meet? please set up the equation as well, the steps.
Answer by lwsshak3(11628)

chasity and megan are riding their bicycles to raise money for the Catholic Charities childrens home. they are 45 miles apart when they begin cycling towards one another and they both start cycling at the same time.if Chasity travels at 14 miles per hour and Megan travels 16 miles per hour, how long will take them to meet? please set up the equation as well, the steps.
***
chasity and megan are moving toward each at the sum of their respective speeds=14+16=30 mph
Travel time=distance/speed=45/30=3/2=1.5
how long will it take for chasity and megan to meet? 1.5 hrs